TheSmartOne (thesmartone):

Does the function has a zero in the given interval? Explain why. This is a question on IVT Intermediate value theorem

TheSmartOne (thesmartone):

f(x) = tan(2x) - 4 on \([-\pi, \pi]\) \(\Large f(-\pi) = tan(2( -\pi)) = \frac{sin(-2\pi)}{cos(-2\pi)}=\frac{-sin(2\pi)}{cos(2\pi)} =0\) \(\Large f(\pi) = tan(2( \pi)) = \frac{sin(2\pi)}{cos(2\pi)} =0\)

OpenStudy (rayep12):

lol there is no like picture?

OpenStudy (rayep12):

nvm

TheSmartOne (thesmartone):

f(x) = tan(2x) - 4 on \([-\pi, \pi]\) \(\large f(-\pi) = tan(2( -\pi))-4 = \frac{sin(-2\pi)}{cos(-2\pi)}-4=0 - 4 =-4\) \(\large f(\pi) = tan(2( \pi))-4 = \frac{sin(2\pi)}{cos(2\pi)} -4=0 - 4 = -4\)

zepdrix (zepdrix):

The function does have zeroes in the given interval (you can graph it to see the zeroes). But the Intermediate Value Theorem can not be applied since the function is not continuous on the given interval. So the IVT can not guarantee us that a zero exists in the interval.

TheSmartOne (thesmartone):

So by IVT, since a and b are equal, the function isn't continuous and we won't know if there is a 0 in between the range unless we use a smaller range that is continuous

zepdrix (zepdrix):

Oh, it fails for two different reasons, yes good call :) We don't have a<0 and b>0 like you said.

TheSmartOne (thesmartone):

Yayay, *high five* We got the same solution at the same time xDD And I did graph it haha https://www.desmos.com/calculator/qoapaubntx

zepdrix (zepdrix):

Is yours graphed properly? Did I do mine incorrectly? https://www.desmos.com/calculator/4y1wt1oras Hmm

zepdrix (zepdrix):

Oh I see, you skewed the y-axis.

zepdrix (zepdrix):

I didn't know you can do the notation like that :d interesting

TheSmartOne (thesmartone):

Yeah, my y-axis has a huge range apparently xD

TheSmartOne (thesmartone):

And likewise, I find yours much more cleaner to use I had to search it up on desmos on how to add a restriction and that's the way they showed so I've been using that for the past week :3

TheSmartOne (thesmartone):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TheSmartOne So by IVT, since a and b are equal, the function isn't continuous and we won't know if there is a 0 in between the range unless we use a smaller range that is continuous \(\color{#0cbb34}{\text{End of Quote}}\) want to correct a little mistake I made it isn't continuous because f(a) and f(b) are equal